### v1

# Triaxial Magnetars as Sources of Fast Radio Bursts

Abstract
I suggest that some of the mysterious temporal properties of Fast Radio Bursts (FRB) may be explained if they are produced by dynamically triaxial magnetars. If the bursts are narrowly collimated along open field lines, then observed repeating FRB may be those in which the moment of inertia tensor is only slightly triaxial and the rotation axis, open field lines and radiation point nearly to the observer. Apparently non-repeating FRB may be triaxial with the direction of open field lines and radiation wandering across the sky, reducing their duty factors by several orders of magnitude. A slightly triaxial moment tensor in repeaters moves the line of sight into or out of the radiation pattern or within it, explaining periods of greater or lesser (or absent) activity, and making the probability of detecting a burst and hence the burst rate vary aperiodically. The dynamics of triaxial bodies may also thwart the coherent integration of gravitational signals from fast-rotating accreting neutron stars.
keywords:
Radio transient sources (2008); neutron stars (1108); magnetars (992)
1 Introduction
Many properties of FRB are mysterious, not least why they emit high brightness radio bursts, a mystery they share with radio pulsars. Here I address three other mysteries within the popular and plausible magnetar model (Popov & Postnov, 2007, 2013; Thornton et al., 2013; Lyubarsky, 2014; Pen & Connor, 2015; Katz, 2016; Spitler et al., 2016) of FRB:
  1. 1.
    Why do the duty factors of repeating FRB and apparently non-repeating FRB differ by several orders of magnitude?
  2. 2.
    Why does the activity of repeating FRB vary by large factors between episodes of intense activity and episodes of apparent inactivity?
  3. 3.
    If the magnetar model of FRB is correct, why have their rotational periods not been found, unlike those of known magnetar Soft Gamma Repeaters (SGR) and Anomalous X-ray Pulsars (AXP, which are SGR between outbursts), whose 2–12 s rotational periods appear as large amplitude modulation of their emissions?
In this paper I attribute these phenomena to a triaxial moment of inertia tensor of a magnetically and elastically (by stress frozen-in to the solid component) distorted neutron star. These differences may be enhanced if magnetar-FRB have larger magnetic fields than the known Galactic magnetar-SGR/AXP, implying larger triaxiality of magnetar-FRB moments of inertia.
2 How Triaxiality May Explain the Three Mysteries?
FRB may be produced by charge bunches moving relativistically along magnetic field lines directed toward the observer, likely from their neutron stars’ polar caps (Kumar, Lu & Bhattacharya, 2017; Beniamini & Kumar, 2025), and beamed like all emission from relativistically moving charges. The observed radiation is sensitive to the orientation of the neutron star and its frozen-in field; it requires alignment of the polar cap open magnetic field lines with the line of sight. By itself, this does not explain the absence of periodicity in FRB, for radio pulsar emission is also narrowly beamed yet its periodicity is its most striking feature.
We define a triaxial moment of inertia tensor as one in which the three differences of principal components $|I_{1}-I_{2}|$, $|I_{3}-I_{1}|$ and $|I_{3}-I_{2}|$ are of comparable magnitudes. For a plausible neutron star these differences are $\ll I_{i}$ for $i=1,2,3$; that inequality affects the time scale on which the magnetic pole (and radiation direction) wanders, but not the fact nor the range of its wandering.
If a neutron star’s moment of inertia is triaxial by this definition then its rotation is not periodic, but more complex. A triaxial moment of inertia tensor results from frozen-in magnetic stress $B^{2}/8\pi=4\times 10^{28}B_{15}^{2}$ dyne/cm2 and frozen-in elastic stress ${\cal O}(10^{29}$ dyne/cm2) (Horowitz & Kadau, 2009). These may be of similar magnitude and each makes the moment of inertia tensor triaxial. Rotational distortion relaxes to axisymmetry on a hydrodynamic or viscous time scale, is likely smaller than the other distortions for rotational periods $\gtrsim 1\,$s, and does not contribute to rigid body triaxiality.
Proposed explanations:
  1. 1.
    Luo et al. (2025) suggested that in repeating FRB the magnetic and spin axes are aligned, while Kumar, Lu & Bhattacharya (2017); Beniamini & Kumar (2025) suggested that FRB radiation is emitted along open field lines. This leads to a possible explanation (Katz, 2025a) of the difference between repeating and apparently non-repeating FRB: Apparently non-repeating FRB have triaxial moment tensors, so that their magnetic poles and directions of radiation wander pseudo-randomly across much of the sky. FRB that are are only minimally tri-axial have their magnetic and rotational axes aligned (dissipation reduces them to their lowest energy state with oblate spheriodal intertia tensors with magnetic, emission and rotation axes aligned); if fortuitously these point to the observer we see them as repeating FRB, while otherwise they are not observable at all.
    If the outbursts that produce the observed FRB are stochastic with intrinsic duty factor DF, that may be ${\cal O}(10^{-5})$ for repeating FRB, then the observed duty factor of a triaxial FRB may be ${\cal O}(\text{DF}\Omega/4\pi\sim\text{DF}/4\pi\gamma^{2})$, where $\Omega$ is the solid angle of emission of the FRB and $\gamma$ the corresponding Lorentz factor of the emitting charge bunches. Triaxiality may explain the large gap between the upper bounds on duty factors of apparently non-repeating FRB and the measured duty factors DF of repeaters as curvature radiation models indicate $\gamma^{2}\sim 10^{5}$.
  2. 2.
    The activity of all repeating FRB varies aperiodically on time scales from hours to months or years; this is distinct from the 16.35 d periodic modulation of activity observed in FRB 180916B (CHIME/FRB Collaboration, 2020) and the $\sim 157$ d periodic modulation of activity observed in FRB 20121102 (Rajwade et al., 2020; Cruces et al., 2021). Periods of intense activity, with bursts detected at rates of nearly 1000/h (Zhang et al., 2025), are interspersed among periods in which few or no bursts are detected.
  3. 3.
    Triaxiality may also explain the failure to find a rotational period in any repeating FRB. We suggest that the magnetic and spin axes and the line of sight are closely, but not exactly, aligned in repeating FRB sources (Beniamini & Kumar, 2025; Luo et al., 2025). Then the likelihood of detecting a burst depends on the angle between the magnetic (open field line) axis and the line of sight. Because the rotation of a triaxial body is not periodic the probability of observing a burst is then not a periodic function of time.
The first two of these phenomena might be attributed to “plasma weather” (intrinsic stochasticity of emission) in a magnetar’s magnetosphere, but with no explanatory or predictive theory other hypotheses should be considered. “Plasma weather” cannot explain the absence of periodicity (Katz, 2025b) in a rotating neutron star with a spheroidal (axisymmetric) moment of inertia tensor because the observable “weather” would be modulated at the rotation period (as is Solar activity).
Triaxiality also implies that while we see only a tiny fraction of the bursts emitted by apparent non-repeaters, we see only a similarly tiny fraction of the repeaters.
In contrast, SGR emit thermal X-rays and soft gamma-rays from the photosphere of an equilibrium pair plasma, radiating into at least $2\pi$ sterad ($4\pi$ sterad if the pair plasma fills the magnetosphere), while AXP emission resembles either SGR emission at much lower intensity or thermal emission from the neutron star surface. These radiate very broadly, but not isotropically, so their emission and ${\cal O}(1)$ rotational modulation is observable from most directions. It is insensitive to the neutron star’s orientation. If SGR/AXP are triaxial, this would not produce behavior like that observed for FRB.
3 Euler Equations
The Euler equations of free rigid body rotation are
[…]
To make order of magnitude estimates, replace the time derivatives by the algebraic factor $1/\tau$, where $\tau$ is a characteristic time scale, the fractions by $\Delta I/I$, a measure of the asymmetry of the moment of inertia tensor, and assume $I_{1}\approx I_{2}\approx I_{3}$ to high accuracy, as the case for a nearly spherical body, but that all the $|I_{i}-I_{j}|$ ($i\neq j$) are of the same order of magnitude.
For a magnetized neutron star with a magnetic field of no special symmetry, a rough estimate is
${\Delta I\over I}\sim 10^{-6}B_{15}^{2}.$ (2)
Frozen elastic anisotropy may provide a similar value of $\Delta I/I$. If all the $\Delta I$ and all the $\omega_{i}$ are comparable in magnitude, then
$\tau\sim{I\over\Delta I}{1\over\omega}.$ (3)
For a young magnetar with $\omega\sim 10\,$/s (0.6 s period) and $\Delta I/I\sim 10^{-6}$ the radiation pattern reorients over the sky on a time scale $\tau\sim 1\,$day. Assuming an intrinsic (but stochastic) burst interval $\eta\sim 10^{2}\,$s, as typically observed for active repeaters (that we assume are stably aligned with the line of sight), this is sufficient to explain the failure to see repetitions from apparent non-repeaters: If FRB beams have a width $\sim 0.003\,$rad, as might be expected if the radiating charges have Lorentz factors $\gamma\sim 300$, the radiating beam moves out of the line of sight in a few minutes. The implied recurrence time of apparent non-repeaters is then ${\cal O}(4\pi\gamma^{2}\eta\sim 3\,$y) if their beam has solid angle $1/\gamma^{2}\sim 10^{-5}\,$ sterad and their wandering is stochastic over all directions.
The mechanics of a triaxial rotor may also explain the failure to find periodicity in active repeaters. The alignment of the beam axis with the direction to the observer is unlikely to be exact (and misalignment may be excited by starquake or outburst activity). If the misalignment is a few times larger than the beam width, then the dynamics of a triaxial rotor shift the beam in and out of the line of sight aperiodically, explaining the periods of higher and lesser or no activity in repeating FRB.
If the misalignment is comparable to the beam width then the probability of detecting a burst depends on how close to the line of sight is the beam axis because the radiated intensity decreases with increasing deviation. This varies aperiodically if the deviations from an isotropic moment tensor are triaxial (comparable in magnitude, however small). If the observer remains close to the beam axis the probability of detecting a burst depends on how far he is from the axis. This may explain the failure to find periodicity in the rate of detection of bursts in active repeating FRB Katz (2025b).
Eq. 3 may also be applied to spinning accreting neutron stars that have been suggested as sources of detectable gravitational waves. In order to emit gravitational waves they must not be axisymmetric. That need not, in principle, require them to be triaxial; for example, a dynamically prolate spheroid spinning about one of its major axes has an oscillating and radiating gravitational quadrupole moment. However, that geometry is contrived and implausible; such a neutron star would relax to a nearly oblate configuration. If spheroidal, it would not radiate gravitational waves; the deviations (resulting from magnetic or elastic stress) from an oblate spheroid would likely make it triaxial as well as having a rotating gravitational quadrupole. For a fast-spinning neutron star with the previously estimated elastic stress $\tau$ might be of order minutes. This would thwart the coherent integration required to detect a weak periodic gravitational strain.
4 Discussion
We have argued that three major puzzles about FRB, the orders-of-magnitude difference between duty factors of repeaters and apparent non-repeaters, the aperiodic variation in activity of repeaters, and the failure to find periodic rotational modulation of the burst rate in active repeaters, may be explained in the neutron stars’ moment of inertial tensors are triaxial (specifically, that all components of the deviation from isotropy are comparable—the tensor is not close to spheroidal). If triaxiality successfully explains these puzzles that would support the model of FRB emission along open field lines from magnetars’ polar caps.
The magnetic fields of the very young (10–100 y old) magnetars suggested (Niu et al., 2025) as sources of FRB may be significantly larger than the $10^{14}$$10^{15}$ Gauss observed in Galactic magnetar-SGR inferred to be $10^{3}$$10^{4}$ years old on the basis of their presence in supernova remnants of that age. This would increase the triaxiality of their moment of inertia tensors and decrease the re-orientation time $\tau$ from the values estimated the parameters of the older Galactic magnetars.
Dissipation will tend to align the axis of greatest moment of inertia with the (conserved) direction of angular momentum. If effective, this implies that apparent non-repeaters, with significant triaxiality, are younger than repeaters, whose radiation direction is nearly constant and either directed to the observer or away from the observer and undetected. Dissipation is offset by excitation by starquakes and SGR outbursts and it is unclear if dissipation dominates, or on what time scales.
Data Availability
This theoretical study produced no new data.
Acknowledgments
I thank T. Piran for discussions and the Hebrew University, Jerusalem, Israel for hospitality during this work.